OpenStudy (anonymous):

what single discount is equivalent to two successive discounts of 15% and 7% ?

OpenStudy (campbell_st):

test it out 15% of $100 = $15 then 7% of 85 = $5.95 the successive discounts are the same as a discount of 20.95%

OpenStudy (anonymous):

There is another simple technique u want to calculate the single discount rate for two successive discount so u can use this formula \[(15 + 7) - \left(\begin{matrix}15 \times 7 \ \over \ 100\end{matrix}\right) = 22 - 1.05 = 20.95\]

jhonyy9 (jhonyy9):

- so i think will be 15% + 7% =22%